OpenStudy (anonymous):

Need help. Do not understand how to work this at all. Simplify the expression. Assume all variables are positive. (k^1/6)^-8/(k^11)^1/3 (is fraction)

OpenStudy (anonymous):

book keeping is all it is

OpenStudy (anonymous):

\[\frac{1}{6}\times (-8)-11\times (\frac{1}{3})\] will be your exponent

OpenStudy (anonymous):

the answer should look like k^?=1/k^?

OpenStudy (anonymous):

so would it be k^-8= 1/k^8